I have no idea where to start on this problem. PLease help.
7x/8-4=x

Answer by midwood_trail(310)

I was unsure how to write it in the first place, I'm sorry.
Its (7x/8) - 4=x The 7/8 is a fraction with the x on top with 7.
There is no need to say sorry. A lot of people do not know how to type math questions using their keyboard. I myself made a few errors when I started about 8 years ago.
Here is your question:
(7x/8) - 4 = x
Multiply each term on both sides of the fractional equation by 8 to remove the fraction given on the left side.
After doing so, we get:
7x - 32 = 8x....This is no longer a fractional equation. It is now called a linear equation. The fraction is gone!
We now solve for x.
-32 = 8x - 7x
-32 = x
===================
How do we know that x = -32 is right?
We plug that answer for x back into the original equation and simplify. If we get the same answer on both sides of the equation, then we will know that x = -32.
====================
THE PROVE:
7(-32)/8 - 4 = -32
-224/8 - 4 = -32
-28 - 4 = -32
-32 = -32...IT CHECKS!!!
Final answer: x = -32
I hope this helps.
Happy Thanksgiving.